# Conditional Denoising Meets Polynomial Modeling: A Flexible Decoupled Framework for Time Series Forecasting

### v6

@@ Abstract
Time series forecasting models are becoming increasingly prevalent due to their critical role in decision-making across various domains. However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between their specific components. In particular, fluctuating patterns and smooth trends within time series exhibit distinct characteristics. In this work, to model complicated temporal patterns, we propose a Conditional Denoising Polynomial Modeling (CDPM) framework, where probabilistic diffusion models and deterministic linear models are trained end-to-end. Instead of modeling the coupled time series, CDPM decomposes it into trend and seasonal components for modeling them separately. To capture the fluctuating seasonal component, we employ a probabilistic diffusion model based on statistical properties from the historical window. For the smooth trend component, a module is proposed to enhance linear models by incorporating historical dependencies, thereby preserving underlying trends and mitigating noise distortion. Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models. [^2] Our code is available at https://github.com/zjt-gpu/CDPM.
@@ 1 Introduction
Diffusion-based methods Lee et al. (2024) have shown significant potential in time series forecasting. We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fails to account for the distinct dynamics of different patterns. Specifically, noise is uniformly added across the entire time series, and denoising is then applied indiscriminately, treating all temporal components equally. However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require different treatment. Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties. This strategy of uniform noise addition and denoising fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting. As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing accuracy.
Typically, a single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s overall forecasting ability. Diffusion models, which excel at modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series. In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends. Therefore, by strategically leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more fully capture the dynamic variations and enhance the final forecasting results.
@@ 4.2 Temporal Feature Structuring
Similarly, we compute $X_{T,t}^{0}$ and $X_{T,s}^{0}$ for predicted window. Decomposition operation isolates the distinct characteristics of time series.
@@ 4.3 Conditional Denoising Seasonal Module
Modeling fluctuating seasonal components in time series is challenging due to their non-stationarity and intricate patterns. We propose CDSM, based on a denoising diffusion model, a popular paradigm in probabilistic modeling. By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
@@
where, for each patch $P_{L,i}$, $\mu_{L,i}$ represents the local mean and $\sigma_{L,i}^{2}$ denotes the variance. The historical data is divided into patches $P$ of length $H$, with each patch capturing localized seasonal signals.
@@ 4.4 Polynomial Trend Module
Different trends exhibit distinct patterns of progression, including periods of growth and decline, making it essential to capture a broader range of trend behaviors. By leveraging Cover’s theorem Cover (1965), we project smooth time series patterns into a polynomial space, facilitating the extraction of both linear and nonlinear trends. Based on this, polynomial modeling Liu et al. (2021) is employed to better capture these diverse trend variations. To fully harness polynomial modeling, we introduce a square root transformation as a nonlinear transformation within the polynomial framework. The square root transformation effectively smooths extreme values, enhancing the framework’s flexibility in capturing a wide range of trend patterns.
@@ 4.5 Optimization
During the training and inference phases, after the trend component $\hat{x}_{T,t}^{0}$ is predicted by the PTM and the seasonal component $\hat{x}_{T,s}^{0}$ is recovered by the CDSM.
The final prediction $\hat{x}_{T}^{0}$ is calculated by integrating the trend and seasonal components as follows:
@@
5 Experiments
5.1 Experimental Settings
Table 1: Overview of the dataset characteristics. Here, $T$ denotes the target series used for prediction.
@@ 5.2 Experiment Results
The experimental results in Table 2 validate the effectiveness of the CDPM for time series forecasting. CDPM’s decoupling strategy, which separates trend and seasonal components, allows for more accurate modeling of temporal patterns. It performs well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise. CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics. On high-dimensional datasets like Electricity and Weather, CDPM maintains superior accuracy. Unlike CSDI and Diffusion-TS, which struggle with long forecast horizons, CDPM handles long-term dependencies effectively. On volatile datasets like Wind, CDPM captures sharp fluctuations and achieves competitive results, though further refinement could improve performance on highly noisy data. Overall, CDPM shows strong generalization across various datasets. In summary, CDPM’s decoupling strategy enhances performance, confirming the theoretical hypothesis that it improves accuracy and interpretability over traditional and diffusion-based models.
@@
Table 3: Ablation study results on ETTh2 and ETTm2.
@@
Removing the condition information injection module reduces performance, with a larger impact on ETTm2, where external factors influence intricate temporal patterns. Replacing the PTM with a linear layer causes a significant performance drop, particularly on ETTm2, underscoring the need for non-linear trend modeling. PTM is crucial for capturing long-term trends and periodic behaviors. Eliminating instance normalization leads to the largest performance decline, especially on ETTm2, highlighting its role in stabilizing training for high-resolution datasets. On ETTh2, instance normalization slightly improves performance, indicating its lesser importance when trends and seasonality are clearer. Removing both the condition information injection module and PTM further degrades performance, emphasizing the necessity of both components.
Evaluation of the Decoupling Strategy.
The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with our framework CDPM outperforming all ablated models. The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling. On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability. Applying our framework directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of our framework, highlighting the importance of proper module assignment. These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures. For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but our framework consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
Table 4: Decoupling validation study results on ETTh2 and ETTm2.
@@
To compare the performance of the specially designed components, PTM and CDSM, in modeling the trend and seasonal components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168. The results highlight the effectiveness of tailored modeling strategies for each component and the benefits of decoupled modeling in improving forecasting accuracy. As shown in Figure 4 (a), PTM excels in capturing intrinsic and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series. In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality. For the seasonal component, which exhibits complex fluctuations, the seasonal module demonstrates superior performance, as shown in Figure 4 (b). CDSM effectively captures sharp seasonal changes, while PTM focuses on underlying regularities in the fluctuations, though it may sometimes miss rapid shifts. These results validate each module’s modeling capabilities for its designated component.
@@
To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon on model performance, we conducted a hyperparameter sensitivity analysis using the ETTh1 dataset. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, while larger values (e.g., $d_{\text{model}}=512$) offer diminishing returns and risk overfitting, highlighting the trade-off between model capacity and efficiency. Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global patterns, whereas larger values introduce redundancy and reduce efficiency. Finally, our framework performs consistently across prediction horizons, with the trend module capturing long-term structure and the seasonal module adapting to fluctuations, ensuring reliable forecasts. Overall, the analysis confirms the balance between complexity and scalability in our framework, making it suitable for diverse forecasting tasks.

### v7

@@ Abstract
Time series forecasting models are becoming increasingly prevalent due to their critical role in decision-making across various domains. However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between their specific components. In particular, fluctuating patterns and smooth trends within time series exhibit distinct characteristics. In this work, to model complicated temporal patterns, we propose a Conditional Denoising Polynomial Modeling (CDPM) framework, where probabilistic diffusion models and deterministic linear models are trained end-to-end. Instead of modeling the coupled time series, CDPM decomposes it into trend and seasonal components for modeling them separately. To capture the fluctuating seasonal component, we employ a probabilistic diffusion model based on statistical properties from the historical window. For the smooth trend component, a module is proposed to enhance linear models by incorporating historical dependencies, thereby preserving underlying trends and mitigating noise distortion. Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models. Our code is available at https://github.com/zjt-gpu/CDPM.
@@ 1 Introduction
Diffusion-based methods Lee et al. (2024) have shown great promise in time series forecasting. However, we argue that their naive application of noise addition and denoising overlooks the heterogeneous nature of temporal dynamics. In particular, these methods uniformly add noise across the entire sequence and apply denoising without distinguishing between different temporal components. However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require different treatment. Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties. These distinct patterns require differentiated treatment. Such indiscriminate processing can distort essential features, especially intrinsic trends and fluctuations, thereby degrading forecasting performance. As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing accuracy.
Typically, a single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s overall forecasting ability. Diffusion models, which excel at modeling complex dynamic processes, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series. In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends. Therefore, by strategically leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more fully capture the dynamic variations and enhance the final forecasting results.
@@ 4.2 Temporal Feature Structuring
Similarly, we compute $X_{T,t}^{0}$ and $X_{T,s}^{0}$. Decomposition operation isolates the distinct characteristics of time series.
@@ 4.3 Conditional Denoising Seasonal Module
Modeling fluctuating seasonal components in time series is challenging due to their intricate patterns. We propose CDSM, based on a denoising diffusion model. By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
@@
where, for each patch $P_{L,i}$, $\mu_{L,i}$ represents the local mean and $\sigma_{L,i}^{2}$ denotes the variance. The historical data is divided into patches $P$ of length $H$.
@@ 4.4 Polynomial Trend Module
Different trends follow distinct trajectories, including phases of growth and decline, highlighting the need to capture diverse trend behaviors. By leveraging Cover’s theorem Cover (1965), we project smooth time series patterns into a polynomial space, facilitating the extraction of both linear and nonlinear trends. Based on this, polynomial modeling Liu et al. (2021) is employed to better capture these diverse trend variations. To fully harness polynomial modeling, we introduce a square root transformation to smooth series and enhance the framework’s flexibility in capturing diverse trend patterns.
@@ 4.5 Optimization
The final prediction $\hat{x}_{T}^{0}$ is calculated by integrating the trend $\hat{x}_{T,t}^{0}$ and seasonal components $\hat{x}_{T,s}^{0}$ as follows:
@@
Table 1: Overview of the dataset characteristics.
5 Experiments
5.1 Experimental Settings
@@ 5.2 Experiment Results
The experimental results in Table 2 validate the effectiveness of the CDPM for time series forecasting. CDPM’s decoupling strategy, which separates trend and seasonal components, allows for more accurate modeling of temporal patterns. It performs well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise. CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics. Unlike CSDI and Diffusion-TS, which struggle with long forecast horizons, CDPM handles long-term dependencies effectively. On volatile datasets like Wind, CDPM captures sharp fluctuations and achieves competitive results, though further refinement could improve performance on highly noisy data. Overall, CDPM shows strong generalization across various datasets.
@@
Table 3: Ablation study results on ETTh2 and ETTm2.
@@
Removing the condition information injection module reduces performance, with a larger impact on ETTm2, where external factors influence intricate temporal patterns. Replacing the PTM with a linear layer causes a significant performance drop, which is crucial for capturing long-term trends and periodic behaviors. Eliminating instance normalization leads to the largest performance decline, highlighting its role in stabilizing training for high-resolution datasets. Removing both the condition information injection module and PTM further degrades performance, emphasizing the necessity of both components.
@@
Table 4: Decoupling validation study results on ETTh2 and ETTm2.
Evaluation of the Decoupling Strategy.
The decoupling strategy that separates trend and seasonal patterns markedly improves performance. Applying our framework directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of our framework, highlighting the importance of proper module assignment. These results emphasize the need to tailor modules for distinct patterns, with decoupling especially beneficial for datasets with clear temporal structure.
@@
To evaluate the effectiveness of PTM and CDSM in modeling different components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168. The results highlight the effectiveness of tailored modeling strategies for each component. As shown in Figure 4 (a), PTM excels in capturing intrinsic and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series. In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality. For the seasonal component, which exhibits complex fluctuations, the seasonal module demonstrates superior performance, as shown in Figure 4 (b). These results validate each module’s modeling capabilities for its designated component.
@@
To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon on model performance, we conducted a hyperparameter sensitivity analysis using the ETTh1 dataset. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, while larger values (e.g., $d_{\text{model}}=512$) offer diminishing returns and risk overfitting, highlighting the trade-off between model capacity and efficiency. Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global patterns, whereas larger values introduce redundancy and reduce efficiency. Finally, our framework performs consistently across prediction horizons.